To hire a trailer the initial cost is $50 plus an additional charge of $10 per hour.
(a) Write an equation for the cost ($C) if you hire the trailer for t hours.

(b) What is the cost if you hire the trailer for 4 hours?

(c) How many hours was the trailer hired for if the total charge was $125?
1
Expert's answer
2020-11-10T12:43:09-0500

a) An equation for the cost ($C) if we hire the trailer for t hours is the following:

"C(t)=50+10t"


b)  The cost if we hire the trailer for 4 hours is "C(4)=50+10\cdot4=50+40=90"


c) If the total charge was $125, then for hours "t" we have the following equation:

"125=50+10t"

which is equivalent to

"10t=125-50"

and to

"10t=75".

And consequently, t = 7.5 hours. Therefore, the trailer was hired 7 hours and 30 minutes.